Engineers want to design seats in commercial aircraft so that they are wide enough to fit 95​%of all males.​ (Accommodating 100% of males would require very wide seats that would be much too​ expensive.) Men have hip breadths that are normally distributed with a mean of 14.7 in. and a standard deviation of1.1in. FindUpper P 95That​ is, find the hip breadth for men that separates the smallest 95​%from the largest55​%.The hip breadth for men that separates the smallest95​%from the largest55​%isUpper P 95= in
Answer #1

Given that,

mean = = 14.7

standard deviation = = 1.1

Using standard normal table,

P(Z < z) = 95%

= P(Z < z) = 0.95  

= P(Z <1.64 ) = 0.95

z = 1.64

Using z-score formula  

x= z * +

x= 1.64*1.1+14.7

x= 16.504

x=17